OpenStudy (anonymous):

14) Juliana has saved her pocket money to buy up to 3 fashion magazines. If there are 5 different magazines to choose from, the number of ways she can buy 1, 2 or 3 magazines is: A) 90 B) 80 C) 25 D) 70 E) 85

OpenStudy (cwrw238):

if she buys 1 its 5 ways 2 - 5C2 = its 5*4/2 = 10 3 its 5c3 = 5*4*3 / 3*2 = 10 that a total of 25

OpenStudy (anonymous):

number of ways she can buy 1 is C(1,5) number of ways she can buy 2 is C(2,5) number of ways she can buy 3 is C(3,5) C(1,5)+C(2,5)+C(3,5)=25

OpenStudy (anonymous):

the answer is 85....

OpenStudy (anonymous):

which is E.

OpenStudy (anonymous):

let's think again

OpenStudy (cwrw238):

ok - if it depends on the order in which she buys the magazines it is 85 - this wasn't mentioned in the question in this case its 5 + 5P2 + 5P3 = 5 + 20 + 60 = 85 - permutations not combinatons

OpenStudy (anonymous):

this question is under permutations so there should be just a simple way to multiply them but I figured out how it is 85 using nPr, 5P1+5P2+5P3=85 ways.

OpenStudy (anonymous):

ur also using the nPr way, how would u do it using just permutations?

OpenStudy (cwrw238):

nPr means number of permutations of r in n and = n! / r!

OpenStudy (anonymous):

hmm, so it's basically the same?

OpenStudy (cwrw238):

yea - oh ill check on that formula - not sure if its right is it correct?

OpenStudy (anonymous):

i think it's nPr=n!/(n-r)!

OpenStudy (cwrw238):

yup - you r right - my memory failed me there and nCr - n! / r! * (n-r)! right?

OpenStudy (cwrw238):

a quick way of calculating say 5C2 is 5*4 --- 2*1 i.e. number of digits in top and bottom are the same

OpenStudy (anonymous):

E.